If area, velocity and density are base units then find the dimensions of force

Grade:11

1 Answers

Koushika
6 years ago
Let us assign the SI units of area ,velocity , density and force to variables
a=m2    
v=ms-1   so, s-1=v/m   ,  s-2=v2/m2
d=kgm-3      so, kg=dm3
force= kg m s-2
           =dm×m×v2/m2
           =dm2v2
          =dav2                (since, m2=a)